# DriveRX: A Vision-Language Reasoning Model for Cross-Task Autonomous Driving

### v1

@@ Abstract
Autonomous driving requires real-time, robust reasoning across perception, prediction, planning, and behavior. However, conventional end-to-end models fail to generalize in complex scenarios due to the lack of structured reasoning. While recent vision-language models (VLMs) have been applied to driving tasks, they typically rely on isolated modules and static supervision, limiting their ability to support multi-stage decision-making. We present AutoDriveRL, a unified training framework that formulates autonomous driving as a structured reasoning process over four core tasks. Each task is independently modeled as a vision-language QA problem and optimized using task-specific reward models, enabling fine-grained reinforcement signals at different reasoning stages. Within this framework, we train DriveRX, a cross-task reasoning VLM designed for real-time decision-making. DriveRX achieves strong performance on the public benchmark, outperforming GPT-4o in behavior reasoning and demonstrating robustness under complex or corrupted driving conditions. Our analysis further highlights the impact of vision encoder design and reward-guided reasoning compression. We will release the AutoDriveRL framework and the DriveRX model to support future research.
@@ 1 Introduction
Autonomous driving systems operate in dynamic and uncertain traffic environments [1, 2, 3], requiring robust, real-time capabilities in perception, prediction, and decision-making [4, 5, 6]. However, in complex scenarios such as occlusion or abnormal behaviors, conventional end-to-end models exhibit limited generalization due to the lack of structured reasoning and semantic interpretability [7, 8, 9].
Recent progress in vision-language models (VLMs) has extended their applications to autonomous driving tasks, including object detection [10, 11, 12], trajectory prediction [13, 14, 15], and planning [16, 15]. Prior work [17, 15, 18] demonstrates VLMs’ potential through language prompts and QA mechanisms, yet primarily focuses on isolated task modules without a unified reasoning framework. This fragmentation hinders systematic understanding and cross-task generalization [19, 18].
However, applying VLMs to autonomous driving remains challenging due to: (1) absence of intermediate reasoning chains [20, 16, 17]; (2) isolation across perception, prediction, planning, and behavior [21, 22, 23]; (3) reliance on static prompts and annotations without adaptive feedback [24, 25, 26].
Meanwhile, large language models (LLMs) have shown promising solutions to similar challenges [27, 28, 29, 30]. OpenAI’s o1 [30] leverages chain-of-thought (CoT) prompting to generate intermediate reasoning steps, while DeepSeek’s R1-Zero [29] uses reinforcement learning (RL) to improve model behavior through dynamic feedback—without relying on supervised fine-tuning. These approaches address key limitations in current VLMs by enabling structured reasoning, unifying task semantics, and supporting feedback-driven optimization, motivating their extension to autonomous driving.
@@
Several multimodal studies have explored integrating reasoning mechanisms to enhance model capabilities. For example, LlamaV-o1 [31] combines Monte Carlo Tree Search (MCTS) with CoT prompting to improve logical consistency in visual question answering (VQA); VLM-R1 [32] and MM-Eureka [33] apply RL and rule-based supervision to extend reasoning into math and science domains. While effective in their domains, these approaches are not designed for the real-time, decision-critical demands of autonomous driving. Long-form reasoning chains, common in CoT-based methods, are often too slow or verbose for driving scenarios that require concise, executable outputs. AlphaDrive [16] explores the application of RL in autonomous driving, but only within the planning task, and does not address multi-task coordination.
This motivates the development of a unified vision-language framework that supports structured reasoning, reinforcement learning, and multi-task coordination across core autonomous driving tasks.
To this end, we propose AutoDriveRL, a multi-task training framework for VLMs in autonomous driving. AutoDriveRL combines language-driven reasoning with task-specific RL strategies, aiming to systematically improve model performance and robustness across core autonomous driving tasks.
@@
AutoDriveRL builds on prior work that decomposes autonomous driving into four core sub-tasks: perception, prediction, planning, and behavior. These tasks are organized into a sequential reasoning chain, but each task is independently modeled as a VQA problem without relying on outputs or contextual information from previous stages. Crucially, AutoDriveRL introduces task-specific reward models that provide fine-grained, stage-level reinforcement signals during training. From a global perspective, this structured supervision enables process-level feedback across reasoning steps, supports interpretable intermediate reasoning, facilitates cross-task optimization, and significantly enhances the model’s stability and overall performance in autonomous driving decision-making.
@@
Within this framework, we train a unified cross-task vision-language model, DriveRX, where “RX” denotes its focus on reasoning and cross-task generalization in autonomous driving. DriveRX consistently outperforms existing models across perception, prediction, planning, and behavior tasks, in both in-distribution and out-of-distribution (OOD) benchmark settings. On the public benchmark DriveBench [34], DriveRX achieves a score of 62.02 in behavior, surpassing GPT-4o’s [27] 55.04 by 6.98 points. It maintains stable outputs under challenging conditions such as rain, occlusion, and lighting variation, demonstrating strong robustness and generalization. On the DriveBench-Corruption subset, DriveRX achieves 55.24, again outperforming GPT-4o (53.97).
Notably, driven by the reward design in AutoDriveRL, DriveRX produces more compact reasoning traces and faster responses than existing reasoning-based models—better aligning with the real-time requirements of autonomous driving.
@@
  • We propose AutoDriveRL, a unified training framework that leverages semantic reward modeling to enable structured reasoning across four core autonomous driving tasks: perception, prediction, planning, and behavior.
  • We develop a cross-task vision-language model, DriveRX, which achieves strong performance across multiple driving tasks and complex traffic scenarios, demonstrating strong generalization and robustness, especially in behavior tasks.
  • We conduct an in-depth analysis of the training dynamics, identifying cross-task interactions, vision encoder bottlenecks in perception, and the impact of reward-guided reasoning compression.
@@ 2.1 Vision-Language Models for Autonomous Driving
Recent works have explored using vision-language models (VLMs) to enhance perception and reasoning in autonomous driving. DriveLM [35] formulates multiple driving tasks as structured graph-based QA, constructing a multi-task dataset. Dolphins [36] focuses on dialogue-based reasoning with historical control and counterfactuals. DriveVLM [15] proposes a three-stage language-driven planning pipeline with templated prompts and task-specific instructions. These methods largely rely on supervised fine-tuning (SFT) over static human-annotated data for task-specific outputs. AlphaDrive [16] introduces reinforcement learning (RL) but is limited to the planning stage, lacking support for end-to-end reasoning and broader task-level optimization. In contrast, we adopt an RL framework that decomposes driving into four semantically distinct subtasks, each equipped with a task-specific reward model to evaluate response quality and guide policy learning. This enables the model to acquire task-aligned reasoning strategies at every stage without static supervision, improving robustness and consistency in complex real-world scenarios.
@@ 2.2 Vision-Language Reasoning Models
Recent years have seen notable progress in VLMs for visual reasoning. Representative works include R1-Onevision [37], which enhances visual-textual reasoning via structured intermediate representations and employs reinforcement learning to optimize reasoning trajectories; MM-Eureka [33], which introduces rule-based “aha” mechanisms for iterative refinement of visual answers; LlamaV-o1 [31], which adopts a step-by-step prompting strategy and achieves strong performance on complex benchmarks ; and VisuoThink [38], which leverages multimodal tree search to emulate human-like slow thinking, achieving strong performance in geometry and spatial reasoning tasks. These models are primarily developed for general-purpose vision-language reasoning tasks, typically under static image and single-turn QA settings. They seldom address complex task structures or action-oriented reasoning objectives, making them less applicable to domains like autonomous driving that require long-horizon, context-dependent reasoning across multiple subtasks.
@@ 3 AutoDriveRL
To enhance the reasoning capability and generalization of vision-language models in complex driving scenarios, we propose AutoDriveRL, a unified training framework. AutoDriveRL decomposes the autonomous driving task into four core sub-tasks—and adopts a consistent question-answering structure across all tasks. Each sub-task is equipped with a task-specific reward function, providing targeted feedback to guide the model toward learning stage-specific reasoning strategies. These components are jointly trained using reinforcement learning to develop a multi-task vision-language model, DriveRX. In this section, we detail the approach across four dimensions: task formulation, data construction, training framework, and reward design. An overview of the overall pipeline is illustrated in Figure 1.
@@ 3.1 Task Formulation
Autonomous driving involves perceiving traffic elements, anticipating agent behaviors, and making timely planning and behavioral decisions in dynamic environments. Following prior work [35, 34], we decompose this process into four sub-tasks—perception, prediction, planning, and behavior—that span the full reasoning chain from low-level perception to high-level decision-making.
@@
  • Behavior: Conduct behavior prediction and strategy selection.
@@ 3.2 Data Construction
We employ a unified scoring and filtering pipeline to control task difficulty, increase semantic diversity, and balance action distributions. Low-quality samples (e.g., with annotation errors or duplicate options) are first removed. For each visual question, we collect response scores $s_{ijkm}$ by sampling $K$ outputs from each of $M$ vision-language models. These scores are computed using task-specific reward models (see prompt in Section I.1), guided by structured prompts to ensure consistency and reproducibility without human labeling. Alignment with human judgments is shown in Appendix F.
@@
Based on $D_{i}$, we follow prior work [29, 39] on RL data filtering strategies and exclude both overly easy and overly hard samples from training. Specifically, we select mid-range frames to improve training stability by avoiding trivial examples and high-variance outliers. Low-scoring frames—typically involving occlusion, poor lighting, or rare behaviors—are used exclusively for evaluation to assess generalization under challenging conditions.
To address action imbalance—where common actions (e.g., going straight) dominate—we apply class balancing by oversampling rare action samples with explicit intent, improving coverage of low-frequency behaviors. Final distributions are shown in Fig. 6.
This process yields a high-quality, difficulty-controlled, and semantically diverse dataset that supports robust multi-task RL and evaluation.
@@ 3.3 Training Framework
Autonomous driving comprises four semantically coupled sub-tasks—perception, prediction, planning, and behavior decision-making—each requiring intermediate reasoning that is hard to supervise directly. Instead of conventional supervised learning, we adopt a reinforcement-learning paradigm in which a single policy is trained across all tasks and updated from reward feedback.
Many tasks, especially planning and behavior prediction, admit multiple plausible outputs (one-to-many mapping). To cope with the resulting high-variance rewards, we optimize the policy with Group Relative Policy Optimization (GRPO) [40]. For every query $q$, the current policy $\pi_{\theta}$ samples $G$ candidate responses $\{o_{1},\dots,o_{G}\}$. The GRPO objective is
@@
where $r_{i,t}=\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\big{/}\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})$ is the token-level policy ratio, $t$ denotes the token position and $\hat{A}_{i,t}$ is the intra-group advantage defined in Section 3.4.
During training, samples from all four tasks are mixed, and the shared parameters are updated jointly. The framework relies solely on rule-based and model-based reward signals, requiring no extra human annotation and thus scaling easily to new tasks and datasets.
@@ 3.4 Reward Model Design
(ii) LLM-based Reward Model. For remaining candidates, we employ an open-source language model to assign a task-aware quality score $R_{i}\in[0,100]$ based on five criteria: correctness, behavioral understanding, contextual relevance, logic and reasoning and clarity of reasoning. We use distinct prompts tailored to each of the four tasks.
The normalized group-level advantage is computed as:
@@
where $\mu_{R}$ and $\sigma_{R}$ are computed over the $G$ responses sampled for the same input. This scalar advantage is shared across tokens and used in the GRPO update in Eq. (2).

### v2

@@ Abstract
Effective autonomous driving hinges on robust reasoning across perception, prediction, planning, and behavior. However, conventional end-to-end models fail to generalize in complex scenarios due to the lack of structured reasoning. While recent vision-language models (VLMs) have been applied to driving tasks, they typically rely on isolated modules and static supervision, limiting their ability to support multi-stage decision-making. We present AutoDriveRL, a unified training framework that formulates autonomous driving as a structured reasoning process over four core tasks. Each task is independently modeled as a vision-language QA problem and optimized using task-specific reward models, enabling fine-grained reinforcement signals at different reasoning stages. Within this framework, we train DriveRX, a cross-task reasoning VLM designed for multi-stage decision-making. DriveRX achieves strong performance on the public benchmark, outperforming GPT-4o in behavior reasoning and demonstrating robustness under complex or corrupted driving conditions. DriveRX serves as a high-level semantic reasoning backbone, producing structured stage-wise reasoning chains that enhance decision consistency. These outputs also provide high-quality supervisory signals for annotation and downstream planning/control models. We release the AutoDriveRL framework and DriveRX to support future research.
@@ 1 Introduction
Autonomous driving systems operate in dynamic and uncertain traffic environments (Galceran et al., 2015; Xu et al., 2014; Marcu et al., 2024), requiring robust capabilities in perception, prediction, and decision-making (Grigorescu et al., 2020; Chen et al., 2015; Janai et al., 2020). However, in complex scenarios such as occlusion or abnormal behaviors, conventional end-to-end models exhibit limited generalization due to the lack of reasoning ability (Chen et al., 2024a; Xu et al., 2024b; Shao et al., 2023). Moreover, EU regulations and academic research emphasize that autonomous driving systems must be interpretable (Atakishiyev et al., 2023; 2024), highlighting the need for models with structured reasoning and semantic understanding.
Vision-Language Models (VLMs) offer a promising way to meet this demand by combining perception and reasoning. Recent progress has extended their applications to autonomous driving, including object detection (Liu et al., 2023; Choudhary et al., 2024; Qian et al., 2024), trajectory prediction (Mao et al., 2023; Nguyen et al., 2024; Tian et al., 2024), and planning (Jiang et al., 2025; Tian et al., 2024). Prior work (Xu et al., 2024a; Tian et al., 2024; Wang et al., 2024) has demonstrated VLMs’ potential through language prompts and QA mechanisms, but most approaches primarily focuse on isolated task modules without a unified reasoning framework. This fragmentation hinders systematic understanding and cross-task generalization (Zhao et al., 2025; Wang et al., 2024).
Building on these observations, existing VLM-based approaches to autonomous driving exhibit several common limitations: (1) absence of intermediate reasoning chains (Nie et al., 2024; Jiang et al., 2025; Xu et al., 2024a); (2) isolation across perception, prediction, planning, and behavior (Hu et al., 2023; Bansal et al., 2018; Casas et al., 2021); (3) reliance on static prompts and annotations without adaptive feedback (Zhang et al., 2024; Cui et al., 2024; Duan et al., 2024).
Meanwhile, Large Language Models (LLMs) have shown promising solutions to similar challenges (OpenAI, 2024a; Team, 2024; Guo et al., 2025; OpenAI, 2024b). OpenAI’s o1 (OpenAI, 2024b) leverages Chain-of-Thought (CoT) prompting to generate intermediate reasoning steps, while DeepSeek’s R1-Zero (Guo et al., 2025) uses Reinforcement Learning (RL) to improve model behavior through dynamic feedback—without relying on supervised fine-tuning. These approaches address key limitations in current VLMs by enabling structured reasoning, unifying task semantics, and supporting feedback-driven optimization, motivating their extension to autonomous driving.
@@
Several multimodal studies have explored integrating reasoning mechanisms to enhance model capabilities. LlamaV-o1 (Thawakar et al., 2025) combines Monte Carlo Tree Search with CoT prompting to improve logical consistency in Visual Question Answering (VQA); VLM-R1 (Shen et al., 2025) and MM-Eureka (Meng et al., 2025) apply RL and rule-based supervision to extend reasoning into science domains. While effective in their domains, these approaches are not designed for the decision-critical demands of autonomous driving. Long-form reasoning chains, common in CoT-based methods, are often exhibit excessive length and latency for driving scenarios that require concise, executable outputs. AlphaDrive (Jiang et al., 2025) explores the application of RL in autonomous driving, but only within the planning task, and does not address multi-task coordination.
This motivates formulating autonomous driving as a structured reasoning process across four stages—perception, prediction, planning, and behavior—and developing a unified vision-language framework for coherent cross-task semantic understanding.
To this end, we introduce AutoDriveRL, which jointly optimizes a unified VLM through task-specific RL signals that enhance each reasoning stage and improve cross-task consistency and robustness.
@@
Building on this four-stage formulation, AutoDriveRL models each stage as an independent VQA problem organized into a sequential reasoning chain. The framework incorporates task-specific reward models that deliver fine-grained, stage-level reinforcement signals, enabling process-level feedback across reasoning steps. This structured supervision supports interpretable intermediate reasoning, facilitates cross-task optimization, and substantially improves the stability and overall performance of autonomous driving decision making.
@@
Within this framework, we train a unified cross-task vision-language model, DriveRX, where “RX” emphasizing reasoning and cross-task generalization in autonomous driving. DriveRX consistently outperforms existing models across perception, prediction, planning, and behavior tasks, in both in-distribution and out-of-distribution (OOD) benchmark settings. On the public benchmark DriveBench (Xie et al., 2025), DriveRX achieves a score of 62.02 in behavior, surpassing GPT-4o’s (OpenAI, 2024a) 55.04 by 6.98 points. It produces stable outputs under corruption-based perturbations (e.g., rain, occlusion, lighting variation), demonstrating strong robustness to visual degradation. DriveRX obtains 55.24 on DriveBench-Corruption, outperforming GPT-4o (53.97), and further shows strong performance on DriveLM-Hard, which focuses on semantically challenging, long-tail scenarios. Together, these results confirm that DriveRX provides stable and interpretable reasoning chains for high-level semantic understanding and consistent cross-task decision making.
Based on DriveRX, we investigate two downstream applications that leverage its high-level reasoning outputs as structured supervisory signals. First, we extend the vocabulary of DriveRX to build the DriveRX-Agent for trajectory prediction. In open-loop testing on nuScenes, DriveRX-Agent surpasses the baseline DriveLM-Agent, indicating that the reasoning chains of DriveRX offer more reliable guidance in complex, dynamic environments. Second, we distill DriveRX’s stage-wise reasoning data into a compact Vision-Language-Action (VLA) model. The reasoning-enhanced supervision substantially improves action prediction quality, enabling the distilled model to achieve competitive closed-loop performance on Bench2Drive. Together, these results validate that DriveRX’s structured reasoning serves as effective teacher signals or reward feedback for downstream planning and control models, benefiting both trajectory prediction and action generation.
@@
  • Propose AutoDriveRL: We propose AutoDriveRL, a unified RL framework that designs task-specific reward models for four core autonomous driving tasks—perception, prediction, planning, and behavior—and jointly optimizes them to enable structured, interpretable reasoning across tasks.
  • Develop DriveRX: We develop DriveRX, a cross-task VLM that demonstrates strong performance across multiple driving tasks and complex traffic scenarios, with strong generalization and robustness, and serving as a reliable backbone for high-level semantic understanding and reasoning-based analysis in autonomous driving.
  • Explore Applications of VLM: We further explore two downstream applications that leverage DriveRX’s structured reasoning outputs as supervisory signals: (i) DriveRX-Agent, which extends the vocabulary for trajectory prediction and achieves competitive open-loop performance; and (ii) DriveRX-VLA, a compact VLA model distilled from DriveRX’s reasoning data, showing strong closed-loop results on Bench2Drive. These applications demonstrate the versatility and effectiveness of reasoning-enhanced VLMs for downstream planning and control.
@@ 2.1 Vision-Language Models for Autonomous Driving
Recent works have explored using vision-language models (VLMs) to enhance perception and reasoning in autonomous driving. DriveLM (Sima et al., 2024) formulates multiple driving tasks as structured graph-based QA, constructing a multi-task dataset. Dolphins (Ma et al., 2024) focuses on dialogue-based reasoning with historical control and counterfactuals. DriveVLM (Tian et al., 2024) proposes a three-stage language-driven planning pipeline with templated prompts and task-specific instructions. These methods largely rely on supervised fine-tuning (SFT) over static human-annotated data for task-specific outputs. AlphaDrive (Jiang et al., 2025) introduces reinforcement learning (RL) but is limited to the planning stage, lacking support for end-to-end reasoning and broader task-level optimization. In contrast, we adopt an RL framework that decomposes driving into four semantically distinct subtasks, each equipped with task-specific rewards to evaluate response quality and guide policy learning. This enables the model to acquire task-aligned reasoning strategies at every stage without static supervision, improving robustness and consistency in complex real-world scenarios.
@@ 2.2 Vision-Language Reasoning Models
Recent years have seen notable progress in VLMs for visual reasoning. Representative works include R1-Onevision (Yang et al., 2025b), which enhances visual-textual reasoning via structured intermediate representations and employs reinforcement learning to optimize reasoning trajectories; MM-Eureka (Meng et al., 2025), which introduces rule-based “aha” mechanisms for iterative refinement; LlamaV-o1 (Thawakar et al., 2025), which adopts step-by-step prompting and achieves strong performance; and VisuoThink (Wang et al., 2025), which leverages multimodal tree search to emulate human-like slow thinking, achieving strong performance in geometry and spatial reasoning tasks. These models are primarily developed for general-purpose vision-language reasoning, typically under static image and single-turn QA settings. They seldom address complex task structures or action-oriented reasoning, making them less applicable to domains like autonomous driving that require long-horizon, context-dependent reasoning across multiple subtasks.
@@ 3 AutoDriveRL
To enhance the reasoning capability and generalization of VLMs in complex driving scenarios, we propose AutoDriveRL, a unified training framework. AutoDriveRL decomposes the autonomous driving task into four core subtasks and adopts a consistent question-answering structure across all tasks. Each sub-task is equipped with a task-specific reward function, providing targeted feedback to guide the model toward learning stage-specific reasoning strategies. These components are jointly trained using reinforcement learning to develop a multi-task vision-language model, DriveRX. In this section, we detail the approach across four dimensions: task formulation, data construction, training framework, and reward design. An overview of the overall pipeline is illustrated in Figure 1.
@@ 3.1 Task Formulation
Autonomous driving involves perceiving traffic elements, anticipating agent behaviors, and making timely planning and behavioral decisions in dynamic environments. Following prior work (Sima et al., 2024; Xie et al., 2025), we decompose this process into four sub-tasks—perception, prediction, planning, and behavior—that span the full reasoning chain from low-level perception to high-level decision-making.
@@
  • Behavior: Conduct behavior prediction and strategy selection. This task provides a discrete summary of the ego vehicle’s final driving decision (e.g., “slowly go straight”).
@@
In addition, the Behavior stage provides a lightweight and interpretable interface between high-level planning intents and downstream action or trajectory generation, yielding a clearer semantic hierarchy and offering stable supervision for subsequent VLA learning.
@@ 3.2 Data Construction
We employ a unified scoring and filtering pipeline to control task difficulty, increase semantic diversity, and balance action distributions. Low-quality samples (e.g., with annotation errors or duplicate options) are first removed. For each visual question, we collect response scores $s_{ijkm}$ by sampling $K$ outputs from each of $M$ VLMs. These scores are computed using task-specific reward models (see prompt in Section M.1), guided by structured prompts to ensure consistency and reproducibility without human labeling. Alignment with human judgments is shown in Appendix I.
@@
Based on $D_{i}$, we follow prior work (Guo et al., 2025; Zeng et al., 2025) on RL data filtering strategies and exclude both overly easy and overly hard samples from training. Specifically, we select mid-range frames to improve training stability by avoiding trivial examples and high-variance outliers. Low-scoring frames—typically involving occlusion, poor lighting, or rare behaviors—are used exclusively for evaluation to assess generalization under challenging conditions.
To mitigate action imbalance, we oversample rare actions, improving coverage of low-frequency behaviors. Final distributions are shown in Figure 6. These process yields a high-quality, difficulty-controlled, and semantically diverse dataset that supports robust multi-task RL and evaluation.
@@ 3.3 Training Framework
Autonomous driving comprises four semantically coupled sub‑tasks—perception, prediction, planning, and behavior decision‑making—each requiring intermediate reasoning that is hard to supervise directly. Instead of conventional supervised learning, we adopt a reinforcement‑learning paradigm in which a single policy is trained across all tasks and updated from reward feedback.
Many tasks, especially planning and behavior prediction, admit multiple plausible outputs (one‑to‑many mapping). To cope with the resulting high‑variance rewards, we optimize the policy with Group Relative Policy Optimization (GRPO) (Shao et al., 2024). For every query $q$, the current policy $\pi_{\theta}$ samples $G$ candidate responses $\{o_{1},\dots,o_{G}\}$. The GRPO objective is
@@
where $r_{i,t}=\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\big/\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})$ is the token‑level policy ratio, $t$ denotes the token position and $\hat{A}_{i,t}$ is the intra‑group advantage defined in Section 3.4.
During training, samples from all four tasks are mixed, and the shared parameters are updated jointly. The framework relies solely on rule‑based and model‑based reward signals, requiring no extra human annotation and thus scaling easily to new tasks and datasets.
@@ 3.4 Reward Model Design
(ii) LLM-based Reward Model. For remaining candidates, we employ an open-source language model to assign a task-aware quality score $R_{i}\in[0,100]$. Crucially, as detailed in Appendix M.1, distinct prompts and evaluation rubrics are tailored to each of the four sub-tasks to address their specific requirements—ranging from coordinate precision in Perception to safety logic in Planning. Despite these domain differences, all tasks utilize this unified scalar range, ensuring that reinforcement signals remain consistent and directly comparable across the heterogeneous tasks. This standardization prevents optimization instability caused by varying score magnitudes and facilitates effective joint training. The normalized group-level advantage is computed as:
@@
where $\mu_{R}$ and $\sigma_{R}$ are computed over the $G$ responses sampled for the same input. This scalar advantage is shared across tokens and used in the GRPO update in Eq. 2.